# Fourier transform pairs and Eisenstein-type series related to Jacobi elliptic functions

### v1

@@ 2 Fourier transform pairs
The infinite series can be evaluated with the help of formula (8), by first replacing $k\mapsto k^{\prime}$ and then $z\mapsto yK/K^{\prime}$:
@@
$I(y)=-2{\textnormal{i}}+\int_{\mathbb{R}+{\textnormal{i}}\epsilon}\frac{\mathrm{cd}(2K^{\prime}u-{\textnormal{i}}K^{\prime},k)}{\sinh(\pi x-\pi{\textnormal{i}}/2)}e^{2\pi{\textnormal{i}}xy}\mathrm{d}x.$
@@
$\begin{split}\frac{\mathrm{sn}(2K^{\prime}x,k)}{\cosh(\pi x)}\leftrightarrow\frac{\mathrm{ns}(2K^{\prime}x,k)}{\sinh(\pi x)},&\quad(19\leftrightarrow 22),\\[2.15277pt] \frac{\mathrm{cn}(2K^{\prime}x,k)}{\cosh(\pi x)}\leftrightarrow\frac{\mathrm{ds}(2K^{\prime}x,k)}{\sinh(\pi x)},&\quad(20\leftrightarrow 23),\\[2.15277pt] \frac{\mathrm{dn}(2K^{\prime}x,k)}{\cosh(\pi x)}\leftrightarrow\frac{\mathrm{cs}(2K^{\prime}x,k)}{\sinh(\pi x)},&\quad(21\leftrightarrow 24).\end{split}$
We will prove identity 19. Using the same technique as in the derivation of identity 13 from identity 1 on page 2, one can then deduce identity 22. The proofs of identities 20, 21, 23, and 24 follow exactly the same steps and are therefore omitted.
Proof of identity 19. For $y\in\mathbb{R}$ and $m\in{\mathbb{N}}$, define
@@
which is equivalent to identity 19, in view of Legendre’s relation for complete elliptic integrals [2, Eq. (19.7.1)]. $\sqcap\kern-8.0pt\hbox{$\sqcup$}$
@@ 3 Eisenstein-type series
The function $\Lambda_{1,1}(s,\tau)$ can be analytically contined to an entire function of $s$ and it satisfies
@@
The functions $\Lambda_{1,2}(s,\tau)$ and $\Lambda_{2,1}(s,\tau)$ can be analytically contined to entire functions of $s$ and they satisfy
@@
$\Lambda_{0,1},\Lambda_{1,0}$ 19 $\Lambda_{0,2},\Lambda_{2,0}$ 21 $\Lambda_{0,3},\Lambda_{3,0}$ 20
@@ Acknowledgements
The research was supported by the Natural Sciences and Engineering Research Council of Canada.

### v2

@@ 2 Fourier transform pairs
The infinite series can be evaluated with the help of formula (8), by first replacing $k\mapsto k^{\prime}$ and then $z\mapsto 2Ky$:
@@
$I(y)=-2{\textnormal{i}}+e^{\pi y}\int_{\mathbb{R}+{\textnormal{i}}\epsilon}\frac{\mathrm{cd}(2K^{\prime}u-{\textnormal{i}}K^{\prime},k)}{\sinh(\pi u-\pi{\textnormal{i}}/2)}e^{2\pi{\textnormal{i}}uy}\mathrm{d}u.$
@@
$\begin{split}\frac{\mathrm{sn}(2K^{\prime}x,k)}{\cosh(\pi x)}\leftrightarrow\frac{\mathrm{ns}(2K^{\prime}x,k)}{\sinh(\pi x)},&\quad(21\leftrightarrow 24),\\[2.15277pt] \frac{\mathrm{cn}(2K^{\prime}x,k)}{\cosh(\pi x)}\leftrightarrow\frac{\mathrm{ds}(2K^{\prime}x,k)}{\sinh(\pi x)},&\quad(19\leftrightarrow 23),\\[2.15277pt] \frac{\mathrm{dn}(2K^{\prime}x,k)}{\cosh(\pi x)}\leftrightarrow\frac{\mathrm{cs}(2K^{\prime}x,k)}{\sinh(\pi x)},&\quad(20\leftrightarrow 22).\end{split}$
We will prove identity 21. Using the same technique as in the derivation of identity 13 from identity 1 on page 2, one can then deduce identity 24. The proofs of identities 19, 20, 22, and 23 follow exactly the same steps and are therefore omitted.
Proof of identity 21. For $y\in\mathbb{R}$ and $m\in{\mathbb{N}}$, define
@@
which is equivalent to identity 21, in view of Legendre’s relation for complete elliptic integrals [2, Eq. (19.7.1)]. $\sqcap\kern-8.0pt\hbox{$\sqcup$}$
@@ 3 Eisenstein-type series
The function $\Lambda_{1,1}(s,\tau)$ can be analytically continued to an entire function of $s$ and it satisfies
@@
The functions $\Lambda_{1,2}(s,\tau)$ and $\Lambda_{2,1}(s,\tau)$ can be analytically continued to entire functions of $s$ and they satisfy
@@
$\Lambda_{0,1},\Lambda_{1,0}$ 21 $\Lambda_{0,2},\Lambda_{2,0}$ 20 $\Lambda_{0,3},\Lambda_{3,0}$ 19
@@ Acknowledgements
The research was supported by the Natural Sciences and Engineering Research Council of Canada. We are grateful to an anonymous referee for carefully reading the paper and for providing very helpful comments.